OpenStudy (anonymous):

(x+1)^2 = ? -3(x+1) = ?

OpenStudy (anonymous):

\[ (x+1)^2 = x^2+2x+1 \] \[ -3(x+1) = -3x - 3 \]

OpenStudy (anonymous):

Thanks!

OpenStudy (lukecrayonz):

Are you saying they're equal to each other?